OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6. y = 1/3x - 8 y = -3x + 12 y = 3x - 24 y = -1/3x - 4 i have other questions as well please help

OpenStudy (anonymous):

Perpendicular Means Opposite Inverse slope, so the Opposite Inverse of 3x is -1/3x. So, plugging that in as the slope of your mystery line, y = mx + b y = -1/3x + b Using the point (6,-6), x is 6 and y is -6. Plugging that into your equation.... -6 = -1/3(6) + b -6 = -2 + b +2 +2 --------------- b = -4 Therefore, y = -1/3x -4

OpenStudy (anonymous):

@chandhooguy d

OpenStudy (anonymous):